# Botender: Supporting Communities in Collaboratively Designing AI Agents through Case-Based Provocations

### v1

@@ 4. Botender
Based on these design goals, we developed Botender, a system that enables the collaborative design of bots in online communities. While Botender’s system architecture is designed to support the creation of diverse AI agents across different community platforms, in this paper, we present the first version of Botender, which supports the design of single-turn, LLM-based conversational bots powered by AI agents for Discord servers. [^7] In this initial version, the AI agents process one user message at a time and generate an appropriate response when needed. In the Discussion (Section 7), we outline specific future directions to expand support for designing AI agents with additional functionalities across different community platforms. In the following subsections, we first walk through Botender’s user interaction workflow, which enables users to collaboratively design their bot by proposing, iterating on, and deploying tasks for the bot to perform within their community platform. We then describe Botender’s case-based provocation algorithm, which generates provocative test cases to encourage user reflection on bot design during the iteration process. Next, we detail the underlying system architecture that supports Botender’s user interaction workflow. Throughout the section, we connect specific features of Botender’s design to the design goals outlined in the previous section, denoted as D1 to D4. Finally, we conclude with implementation details.
@@
Figure 2. Botender’s proposal page. The left navigation bar lets users switch between viewing all active tasks on their Discord servers, community proposals for desired changes, or experimenting with the bot in the playground without affecting their server. In the center, users see the proposal’s title, description, and the latest proposed edits to the bot’s tasks, such as adding a new task in this screenshot. Users can upvote or downvote to indicate their support for or opposition to deploying the latest edit. The bottom displays a full edit history, allowing users to compare edits with previous versions and the original task. On the right, test cases help guide collaborative decision-making. At the bottom are test cases automatically generated to provoke user reflection and discussion around the latest edit. Generated cases are saved if a user chooses to vote on the bot’s response. At the top, members can review and vote on test cases that have previously been saved by community members. Clicking a test case opens a pop-up with case details, including how the bot’s responses for that case have changed across edits. Finally, users can click “enter other cases manually” to open a sheet where they can add custom test cases.:
@@ 4.1. Proposing Desired Changes to Bot Behavior
When creating a new proposal, users are asked to enter a title and a brief, high-level description of their desired changes. They are also encouraged to include a first attempt at operationalizing the change by either editing an existing task or creating a new one (see Figure 2 for an example). As shown in Figure 3, each task consists of a pair of prompt fields: a trigger and an action. The trigger defines when the bot should perform the task, while the action specifies what the bot should do when the task is triggered (Figure 6). [^9] Following best practices in agentic architectures, although there is just one bot named Botender from users’ perspective, behind the scenes the bot is powered by multiple specialized task agents, coordinated by an orchestrator agent. The trigger prompt is used by the orchestrator, and the action prompt is used by a task agent. The task is also given a brief task name for future reference within the web interface and Discord, but this name does not affect the bot’s behavior. As shown in Figure 4, the short name of the triggered task is displayed to users on Discord each time the bot replies. This allows users to better pinpoint and propose more precise edits to a specific task based on their observations of how the bot actually behaves within their community in accordance with that task.
Once submitted, the proposal is created as a page visible to all community members, as shown in Figure 2. If the original proposer has included a specific proposed edit with their proposal, they and any other visitors to the page are immediately shown a set of test cases illustrating how the bot would behave across a range of interaction scenarios (shown on the right side of Figure 2), and are asked to review these test cases for unintended or undesirable bot behaviors. Each test case shows the channel name where a user message is sent, the user message itself, the name of the triggered task, and how the edited bot would reply to the user message. The test cases are divided into two sections: generated test cases and saved test cases. Generated test cases are produced by Botender’s case-based provocation algorithm (presented below in Section 4.4 with the aim of supporting user reflection on potential opportunities to improve the proposed edit (D2) and helping to surface disagreements among community members about whether the bot’s response in a given scenario is appropriate (D1). The saved test cases section allows the proposer or other community members to save test cases for later consideration and discussion. When saving a test case—whether a generated test case or a bot interaction from a user’s manual testing—the user is asked to thumbs up or thumbs down the test case to indicate whether they think it is an example of a good or bad bot response. Other users can subsequently add their own votes to indicate their own perspectives. If a proposal is created based on a specific bot interaction the user observed in the playground, this case is automatically added to the saved test cases when the proposal is created.
@@ 4.2. Iterating on Bot Behavior
Figure 3. After clicking the edit button on the proposal page, the original static text is replaced by this edit interface. Before saving edits, users are required to run “Test + Generate” to see how the bot would behave with their proposed edits.:
@@
Any user can make an edit to a task from a given proposal page, including creating new tasks or editing or removing existing ones. To do this, users click the “Edit” button next to the current proposed edit. To test their edits, users click the “Test + Generate” button, as shown in Figure 3. Botender then re-runs all saved test cases and displays the updated bot responses based on users’ edits, to support regression testing (D3). Botender also generates new case-based provocations based on the user’s edit. Before a user is able to save their edit, they are asked to review the bot’s behavior in the presented test cases, and must vote on at least one of these test cases. This default threshold was chosen to promote user engagement with the test cases before saving and sharing proposed edits, while avoiding introducing too much friction into the collaborative editing process. However, the threshold may be customized to meet individual communities’ needs. If the user notices a potential issue when reflecting on the test cases, they are encouraged to refine their edit and test again before saving. Once they are satisfied, they can save their edits to the proposal.
Once an edit is saved, the system sends a notification to the proposal’s discussion thread on Discord to encourage people to review the latest proposed edit (D1, D4), as shown in Figure 4. On the proposal page, they can then review the latest proposed edit and corresponding test cases. Users can also view the full edit history of a proposal, including edits to tasks and any additions or removals of saved test cases, and can revert changes as needed (D1). This design mirrors systems such as Wikipedia by making the edit history transparent (Dabbish et al., 2014; Stuart et al., 2012), which helps coordinate efforts and ensures accountability among editors.
Figure 4. The Discord interface, highlighting Botender’s integration with the community platform. (1) By default, Botender replies ”hello” to users who greet it in the #botender channel, as defined by its default “Hello Botender” task. (2) When a new proposal is created, the system sends a notification to the #botender channel, and (3) creates an associated discussion thread, as shown on the right, where users can discuss the proposal and receive notifications about saved edit updates. (4) Once a proposal is deployed, the system notifies the discussion thread, closes the proposal, and (5) and sends a message to the main #botender channel. (6) The bot will then behave according to the latest deployed edit.:
@@ 4.4. Case-Based Provocation Algorithm
A key aspect of Botender is its support for collaborative iteration on bot design through the use of concrete test cases. Drawing inspiration from the use of cases as a medium for collaborative design and deliberation in various fields (Section 2.3), Botender’s case-based provocation algorithm is designed to generate cases that encourage user reflection and discussion about desirable bot behavior, rather than simply validating expected outcomes. In the context of community bots, these provocative cases may highlight situations users may not have considered or instances where community members may disagree on how the bot should behave.
@@
Figure 5. Botender’s case-based provocation algorithm uses three parallel LLM pipelines to generate provocative test cases that encourage user reflection and discussion on common prompt design pitfalls, including ambiguous language, overly narrow phrasing, or unintended downstream consequences for the community. Each pipeline includes its own detector, generator, and evaluator to generate relevant cases. Finally, a selector chooses the most provocative cases from all case candidates. The prompts for all ten LLM modules, including each pipeline’s detector, generator, and evaluator, as well as the final selector, are provided in Appendix A.2.:
Figure 6. Botender’s overall system and agent architecture. On the left, platform events are captured by Botender’s always-running event listener, which translates them into information the agent architecture can understand. The orchestrator agent assesses each event and determines which, if any, task-specific agent is most relevant. The selected task-specific agent then generates an action instruction that is executed by Botender’s platform action executor. On the right, Botender’s website serves as the primary interface for users to collaboratively and iteratively design AI agents. This process generates concrete interaction scenarios that help guide further design iterations. The website is tightly integrated with the community platform, with each proposal directly linked to a dedicated discussion thread, encouraging broader community participation and discussion.:
@@
4.5. System and Agent Architectures
Figure 6 depicts Botender’s overall system architecture and its integration with community platforms such as a Discord server. Behind the scenes, the bot that users interact with is powered by multiple LLM-based agents, including an orchestrator agent and several task-specific agents, each corresponding to a task created by the users. This design aligns with best practices in agentic architectures (Fourney et al., 2024; Epperson et al., 2025), allowing each agent to focus on a single task and thereby enhancing overall transparency, controllability, and task outcomes (Wu et al., 2022; Zhang and Arawjo, 2025). Each task-specific agent includes a name, a trigger prompt, and an action prompt, which users can edit from Botender’s web interface.
When the community platform induces an event detected by Botender’s always-running listener, such as a user sending a message in a channel, the orchestrator agent assesses whether the event is relevant to any of the task-specific agents. This assessment is based on each agent’s trigger prompt. If a specific task is deemed relevant, the corresponding agent will take the necessary action according to its action prompt, such as generating a response to the user’s message. This action is then executed on the community platform. The complete system prompts for both the orchestrator agent and the task-specific agents are provided in Appendix A.1.
It is worth noting that this system architecture is designed to support the bot’s functionality beyond the current scope of single-turn conversations. For example, Botender’s listener can monitor additional platform events, such as new users joining a server, and the system can execute a wider range of platform actions, such as banning users from the server. In the Discussion, we discuss future directions for expanding the bot’s capabilities (Section 7).
@@ 4.6. Implementation Details
Botender is a full-stack, end-to-end system that individual communities on Discord can set up via an installation link, as is standard for Discord bots. Botender’s website has a fixed URL where Discord users can log in with their Discord accounts, but they can only view and design bots for servers where they are members and where Botender is installed. Botender is built with SvelteKit and hosted on Vercel, using shadcn-svelte components styled with Tailwind CSS on the frontend and Firestore databases on the backend. Since Vercel is a serverless platform, Botender’s always-on listener for Discord platform events is hosted separately on Railway. All the AI agents and LLM modules are powered by OpenAI’s GPT-4.1. The entire Botender codebase is open source and publicly available on GitHub. [^11] The link has been removed for review.
5. Validation Study
@@ 5.1. Study Procedures
We conducted the validation study through an online survey. Each participant was randomly assigned to review one of nine pre-selected bot prompts, which we referred to as bot “instructions” in the survey to avoid technical jargon. We selected these prompts to cover a variety of potential pitfalls in prompt design. Each prompt included both a trigger and an action, consistent with how tasks are specified in Botender. All prompts used in the validation study are available in Appendix B.2.
@@
Finally, participants selected the set of cases that revealed more potential problems with the prompt and thus better highlighted ways to improve the bot. They also briefly justified their choice.
5.2. Recruitment
We recruited 90 participants on Prolific with experience in online groups or communities (e.g., Discord servers or Slack workspaces) where Botender is intended to be deployed. Each participant was compensated \$5 USD, and the median survey completion time was 13 minutes. This sample size yielded 10 independent reviews per prompt and its corresponding cases.
5.3. Validation Study Results
Figure 7. The number of participants who reported that the sets of cases generated by the baseline vs. Botender’s case-based provocation algorithm were more provocative (i.e., revealed more opportunities to improve the bot’s instructions).:
Overall, participants found that Botender’s case-based provocations revealed more opportunities to improve the bot’s instructions. As shown in Figure 7, most participants selected the set of cases generated by Botender’s algorithm when asked which set revealed more opportunities to improve the bot’s instructions, in a blind comparison. Participants’ justification of their choices [^12] We denote participant IDs starting with V and the prompt a participant reviewed starting with P—for example, “(V79, P1)”. All prompts and cases participants reviewed can be found in Appendix B.2. offered insight into the differences they perceived between the two sets of cases. For example, several participants noted that Botender’s case-based provocations “surface a wider range of edge cases that the current instruction doesn’t handle well” (V80, P8). Participants mentioned that these cases “are more vague and difficult for the bot to interpret” (V79, P1), “show tricky situations the bot’s instructions don’t cover” (V59, P7), and “highlight unclear criteria and over-triggering, thus better exposing instruction weaknesses” (V62, P2). Meanwhile, they found the standard test cases to be “more cut and dry” (V4, P7), offering either “obvious examples of red flags to the bot” (V19, P1) or cases that “shows the bot doing its job correctly” (V72, P2).

### v2

@@ 4. Botender
Based on these design goals, we developed Botender, a system that enables the collaborative design of bots in online communities. While Botender’s system architecture is designed to support the creation of diverse AI agents across different community platforms, in this paper, we present the first version of Botender, which supports the design of single-turn, LLM-based conversational bots powered by AI agents for Discord servers. In the following subsections, we first introduce Botender’s overall system and agent architecture. Next, we walk through Botender’s user interaction workflow, as illustrated in Figure 1. This workflow enables users to collaboratively design their bot by proposing, iterating on, and deploying tasks for the bot to perform within their community platform. We then describe Botender’s case-based provocation algorithm, which generates provocative test cases to encourage user reflection on bot design during the iteration process. Throughout the section, we connect specific features of Botender’s design to the design goals outlined in the previous section, denoted as D1 to D4. Finally, we conclude with implementation details.
4.1. System and Agent Architectures
Figure 2. Botender’s overall system and agent architecture. On the left, platform events are captured by Botender’s always-running event listener, which translates them into information the agent architecture can understand. The orchestrator agent assesses each event and determines which, if any, task-specific agent is most relevant. The selected task-specific agent then generates an action instruction that is executed by Botender’s platform action executor. On the right, Botender’s website serves as the primary interface for users to collaboratively and iteratively design AI agents. This process generates concrete interaction scenarios that help guide further design iterations. The website is tightly integrated with the community platform, with each proposal directly linked to a dedicated discussion thread, as shown in Figure 5. This integration encourages broader community participation and discussion.
An illustration of Botender’s system architecture, showing four main components: community platform, platform integration, community AI agents, and Botender’s web app.
Figure 2 depicts Botender’s overall system architecture and its integration with community platforms such as a Discord server. From the user’s perspective, they install a single bot named Botender and interact with it on their community platform. [^7] Like any Discord bot, only server admins can install Botender on their Discord server. After installation, admins can then grant other users permission to collaborate on designing the bot’s behavior together. They can also change the bot’s name from Botender to any name of their choosing. The bot can perform a variety of tasks. The current version of Botender supports single-turn tasks: the bot handles one user message at a time, along with some meta information like the channel where the message was posted, and responds if any of its available tasks are triggered. For example, as shown on the left side of Figure 2, if a user posts an advertisement in a certain channel, the bot recognizes this as relevant to its “redirect ads” task. It then replies to the user, suggesting they post the ad in a more appropriate channel. Users can work together to create any kind or number of tasks they want the bot to handle, tailoring its behavior to fit the unique needs and preferences of their community.
Behind the scenes, as shown on the right side of Figure 2, the bot that users interact with is powered by multiple LLM-based agents, including an orchestrator agent and several task-specific agents, each corresponding to a task created by the users. This design aligns with best practices in agentic architectures (Fourney et al., 2024; Epperson et al., 2025), allowing each agent to focus on a single task and thereby enhancing overall transparency, controllability, and task outcomes (Wu et al., 2022; Zhang and Arawjo, 2025). Each task-specific agent includes a name, a trigger prompt, and an action prompt, which users can edit from Botender’s web interface. When the community platform induces an event detected by Botender’s always-running listener, such as a user sending a message in a channel, the orchestrator agent assesses whether the event is relevant to any of the task-specific agents. This assessment is based on each agent’s trigger prompt. If a specific task is deemed relevant, the corresponding agent will take the necessary action according to its action prompt, such as generating a response to the user’s message. This action is then executed on the community platform. [^8] Botender’s system architecture is designed to support the bot’s functionality beyond conversational interactions. For example, the event listener can monitor additional platform events, such as new users joining a server, and the system can execute a wider range of platform actions, such as banning users from the server. In the Discussion (Section 7), we outline future directions for expanding the bot’s capabilities. The complete system prompts for both the orchestrator agent and the task-specific agents are provided in Appendix A.1.
In the following subsections, we next walk through how users can use Botender to collaboratively design community bot behavior, by (1) proposing changes, (2) testing and iterating on prompts to operationalize proposed changes, and (3) deploying changes to the community platform (Figure 1).
@@ 4.1. System and Agent Architectures
Figure 3. Botender’s proposal page. The left navigation bar lets users switch between viewing all active tasks on their Discord servers, community proposals for desired changes, or experimenting with the bot in the playground without affecting their server. In the center, users see the proposal’s title, description, and the latest proposed edits to the bot’s tasks, such as adding a new task in this screenshot. Users can upvote or downvote to indicate their support for or opposition to deploying the latest edit. The bottom displays a full edit history, allowing users to compare edits with previous versions and the original task. On the right, test cases help guide collaborative decision-making. At the bottom are test cases automatically generated to provoke user reflection and discussion around the latest edit. Generated cases are saved if a user chooses to vote on the bot’s response. At the top, members can review and vote on test cases that have previously been saved by community members. Clicking a test case opens a pop-up with case details, including how the bot’s responses for that case have changed across edits. Finally, users can click “enter other cases manually” to open a sheet where they can add custom test cases.
A screenshot of Botender’s proposal page.
@@ 4.2. Proposing Desired Changes to Bot Behavior
When creating a new proposal, users are asked to enter a title and a brief, high-level description of their desired changes. They are also encouraged to include a first attempt at operationalizing the change by either editing an existing task or creating a new one (see Figure 3 for an example). As shown in Figure 4, each task consists of a pair of prompt fields: a trigger and an action. The trigger defines when the bot should perform the task, while the action specifies what the bot should do when the task is triggered. The task is also given a brief task name for future reference within the web interface and Discord, but this name does not affect the bot’s behavior. As shown in Figure 5, the short name of the triggered task is displayed to users on Discord each time the bot replies. This allows users to better pinpoint and propose more precise edits to a specific task based on their observations of how the bot actually behaves within their community in accordance with that task. [^10] These name, trigger, and action fields correspond to each task-specific agent’s name, trigger prompt, and action prompt as discussed in Section 4.1.
Once submitted, the proposal is created as a page visible to all community members, as shown in Figure 3. If the original proposer has included a specific proposed edit with their proposal, they and any other visitors to the page are immediately shown a set of test cases illustrating how the bot would behave across a range of interaction scenarios (shown on the right side of Figure 3), and are asked to review these test cases for unintended or undesirable bot behaviors. Each test case shows the channel name where a user message is sent, the user message itself, the name of the triggered task, and how the edited bot would reply to the user message. The test cases are divided into two sections: generated test cases and saved test cases. Generated test cases are produced by Botender’s case-based provocation algorithm (presented below in Section 4.5) with the aim of supporting user reflection on potential opportunities to improve the proposed edit (D2) and helping to surface disagreements among community members about whether the bot’s response in a given scenario is appropriate (D1). The saved test cases section allows the proposer or other community members to save test cases for later consideration and discussion. When saving a test case—whether a generated test case or a bot interaction from a user’s manual testing—the user is asked to thumbs up or thumbs down the test case to indicate whether they think it is an example of a good or bad bot response. Other users can subsequently add their own votes to indicate their own perspectives. If a proposal is created based on a specific bot interaction the user observed in the playground, this case is automatically added to the saved test cases when the proposal is created.
@@ 4.3. Iterating on Bot Behavior
Figure 4. After clicking the edit button on the proposal page, the original static text is replaced by this edit interface. Before saving edits, users are required to run “Test + Generate” to see how the bot would behave with their proposed edits.
An abridged screenshot of the edit interface on Botender’s proposal page.
@@
Any user can make an edit to a task from a given proposal page, including creating new tasks or editing or removing existing ones. To do this, users click the “Edit” button next to the current proposed edit. To test their edits, users click the “Test + Generate” button, as shown in Figure 4. Botender then re-runs all saved test cases and displays the updated bot responses based on users’ edits, to support regression testing (D3). Botender also generates new case-based provocations based on the user’s edit (D2). Before a user is able to save their edit, they are asked to review the bot’s behavior in the presented test cases, and must vote on at least one of these test cases. This default threshold was chosen to promote user engagement with the test cases before saving and sharing proposed edits, while avoiding introducing too much friction into the collaborative editing process. However, the threshold may be customized to meet individual communities’ needs. If the user notices a potential issue when reflecting on the test cases, they are encouraged to refine their edit and test again before saving. Once they are satisfied, they can save their edits to the proposal.
Once an edit is saved, the system sends a notification to the proposal’s discussion thread on Discord to encourage people to review the latest proposed edit (D1, D4), as shown in Figure 5. On the proposal page, they can then review the latest proposed edit and corresponding test cases. Users can also view the full edit history of a proposal, including edits to tasks and any additions or removals of saved test cases, and can revert changes as needed (D1). This design mirrors systems such as Wikipedia by making the edit history transparent (Dabbish et al., 2014; Stuart et al., 2012; Kuo et al., 2024), which helps coordinate efforts and ensures accountability among editors.
Figure 5. The Discord interface, highlighting Botender’s integration with the community platform. (1) By default, Botender replies ”hello” to users who greet it in the #botender channel, as defined by its default “Hello Botender” task. (2) When a new proposal is created, the system sends a notification to the #botender channel, and (3) creates an associated discussion thread, as shown on the right, where users can discuss the proposal and receive notifications about saved edit updates. (4) Once a proposal is deployed, the system notifies the discussion thread, closes the proposal, and (5) and sends a message to the main #botender channel. (6) The bot will then behave according to the latest deployed edit.
The Discord interface displaying Botender’s integration with the community platform, featuring six circular labels that correspond to the integration details described in the figure caption.
@@ 4.5. Case-Based Provocation Algorithm
A key aspect of Botender is its support for collaborative iteration on bot design through the use of concrete test cases. Drawing inspiration from the use of cases as a medium for collaborative design and deliberation in various fields (Section 2.3), Botender’s case-based provocation algorithm is designed to generate concrete interaction scenarios that spark collective reflection and discussion about desirable bot behavior, rather than simply validating expected outcomes. The goal of these case-based provocations is to test the boundaries of users’ intents in bot design (as expressed through LLM prompts) by (1) highlighting potentially undesired bot behavior in interaction scenarios users may not have considered; and (2) highlighting areas where there is potential for disagreement among different users. Based on past research, discussed in Section 2.3, we expect such boundary-challenging provocations to be more useful in supporting prompt iteration than test cases aimed at simply validating expected behavior in expected scenarios.
@@
Figure 6. Botender’s case-based provocation algorithm uses three parallel LLM pipelines to generate provocative test cases that encourage user reflection and discussion on common prompt design pitfalls, including ambiguous language, overly narrow phrasing, or unintended downstream consequences for the community. Each pipeline includes its own detector, generator, and evaluator to generate relevant cases. The detector identifies phrases related to the targeted pitfall; the generator creates cases based on these identified phrases; and the evaluator checks how well the generated cases illustrate the pitfall. Finally, a selector chooses the most provocative cases from all case candidates. The prompts for all ten LLM modules, including each pipeline’s detector, generator, and evaluator, as well as the final selector, are provided in Appendix A.2.
A flowchart illustrating the LLM modules that power Botender’s case-based provocation algorithm.
@@ 4.6. Implementation Details
Botender is a full-stack, end-to-end system that individual communities on Discord can set up via an installation link, as is standard for Discord bots. Botender’s website has a fixed URL where Discord users can log in with their Discord accounts, but they can only view and design bots for servers where they are members and where Botender is installed. Botender is built with SvelteKit and hosted on Vercel, using shadcn-svelte components styled with Tailwind CSS on the frontend and Firestore databases on the backend. Since Vercel is a serverless platform, Botender’s always-on listener for Discord platform events is hosted separately on Railway. All the AI agents and LLM modules are powered by OpenAI’s GPT-4.1. The entire Botender codebase is open source and publicly available on GitHub. [^12] https://github.com/tskuo/Botender
5. Algorithm Validation Study
@@
5.1. Recruitment
We conducted the validation study through an online survey. In total, we recruited 90 participants on Prolific with experience in online groups or communities (e.g., Discord servers or Slack workspaces) where Botender is intended to be deployed. Each participant was compensated \$5 USD, and the median survey completion time was 13 minutes.
@@ 5.2. Study Procedures
Each participant was randomly assigned to review one of nine pre-selected bot prompts, which we referred to as bot “instructions” in the survey to avoid technical jargon. We selected these prompts to cover a variety of potential pitfalls in prompt design. Each prompt included both a trigger and an action, consistent with how tasks are specified in Botender. Additional details about the prompts used in the validation study are available in Appendix B.2.
@@
Finally, participants selected the set of cases that revealed more potential problems with the prompt and thus better highlighted ways to improve the bot. They also briefly justified their choice. With a total of 90 participants, this sample size yielded 10 independent reviews per prompt and its corresponding cases. All prompts and cases that participants reviewed can be found in the supplementary materials.
5.3. Study Results
Figure 7. The number of participants who reported that the sets of cases generated by the baseline vs. Botender’s case-based provocation algorithm were more provocative (i.e., revealed more opportunities to improve the bot’s instructions).
A bar chart comparing the number of participants who found the sets of cases generated by the baseline and by Botender’s case-based provocation algorithm to be more provocative. The results show 17 participants favored the baseline, 15 reported no difference, and 58 preferred Botender.
Overall, participants found that Botender’s case-based provocations revealed more opportunities to improve the bot’s instructions. As shown in Figure 7, most participants selected the set of cases generated by Botender’s algorithm when asked which set revealed more opportunities to improve the bot’s instructions, in a blind comparison. Participants’ justification of their choices [^13] We denote participant IDs starting with V and the prompt a participant reviewed starting with P—for example, “(V79, P1)”. offered insight into the differences they perceived between the two sets of cases. For example, several participants noted that Botender’s case-based provocations “surface a wider range of edge cases that the current instruction doesn’t handle well” (V80, P8). Participants mentioned that these cases “are more vague and difficult for the bot to interpret” (V79, P1), “show tricky situations the bot’s instructions don’t cover” (V59, P7), and “highlight unclear criteria and over-triggering, thus better exposing instruction weaknesses” (V62, P2). Meanwhile, they found the standard test cases to be “more cut and dry” (V4, P7), offering either “obvious examples of red flags to the bot” (V19, P1) or cases that “shows the bot doing its job correctly” (V72, P2).